OpenStudy (anonymous):

Solve for x: 4/5 x + 4/3 = 2x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

Is that (4)/(5x) or (4/5)x?

OpenStudy (anonymous):

(4/5)x

OpenStudy (texaschic101):

There is two ways to do this... (1) 4/5x + 4/3 = 2x --- multiply by LCD, which is 15 12x + 20 = 30x 20 = 30x - 12x 20 = 18x 20/18 = x 10/9 = x (2) 4/5x + 4/3 = 2x -- subtract 4/5x 4/3 = 2x - 4/5x 4/3 = 10/5x - 4/5x 4/3 = 6/5x (4/3) / (6/5) = x 4/3 * 5/6 = x 20/18 = x 10/9 = x you arrive at the same answer...it just depends on if you want to work the equation with fractions in there or not

OpenStudy (anonymous):

the first one I can understand :) thank you very much

OpenStudy (texaschic101):

anytime :)